Suppose there are only two firms in an economy: Cowhide, Inc. produces leather and sells it to Couches, Inc which produces and sells leather furniture. With each $1,000 of leather that it buys from Cowhide, Inc, Couches, Inc. produces a couch and sells it for $3,000. Neither firm had any inventory at the beginning of 2009, During that year, Cowhide produced enough leather for 20 couches. Couches, Inc. bought 80% of that leather for $16,000 and promised to buy the remaining 20% for $4,000 in 2010, Couches, Inc. produced 16 couches during 2009 and sold each one during that year for $3,000. What was the economys GDP for 2009? a. $52,000 b. $68,000 c. $64,000 d. $48,000
Answer #1

D) $48,000

$4,000 worth of inventory to be sold in 2010 will not be counted for GDP calculation of 2009.

GDP addition by Cowhide Inc in 2009= 16*1000=$16,000

GDP addition by Couches Inc in 2009= Value of output-Intermediate consumption=$3,000*16-(16*$1000)=$32,000

Total GDP for 2009= $32,000+ $16,000= $48,000